The temperature in Adelaide on Christmas Day is predicted to be 40 degrees Celsius. Using linear equation find what the temperature will be in degrees Fahrenheit.
Answer by sachi(548)

the relation between C & F is
C/5=(F-32)/9
or F=9C/5+32
so F=9*40/5+32=104